OpenStudy (liliakarina):

(2*5^2)^5/5^10*2^3=??? I don't get it.

OpenStudy (anonymous):

\[\frac{(2\times 5^2)^5}{5^{10}\times 2^3}\]?

OpenStudy (liliakarina):

Yes I know I have to add the exponents, but I get different results from the ones in the options.

OpenStudy (anonymous):

ok first we do it in our heads, and then we write the math the numerator will bet 2 to the 5th times 5 to the 10th the denominator is 2 to the 3rd times 5 to the 10th the \(5^{10}\) cancel, leaving 1 2 to the 5th over 2 to the 3rd is 2 squared is 4 answer is 4 now we write the math

OpenStudy (anonymous):

\[\frac{(2\times 5^2)^5}{5^{10}\times 2^3}\] \[=\frac{2^5\times 5^{2\times 5}}{5^{10}\times 2^3}\] \[=\frac{2^5\times 5^{10}}{5^{10}\times 2^3}\] \[=\frac{2^5\times \cancel{5^{10}}}{2^3\times \cancel{5^{10}}}\] \[=\frac{2^5}{2^3}=2^{5-3}=2^2=4\]

OpenStudy (liliakarina):

I see now. I missed the last step. Thank you @satellite73

OpenStudy (anonymous):

yw